OpenStudy (anonymous):

how do you explain the relationship and the value of sin 0 can be used to find the other trigonometric values

OpenStudy (anonymous):

u can find the sin of ANY angle using the following relation:\[\sin(x) = \sin(x) - \sin(0)\]it's relly useful

OpenStudy (anonymous):

so basicaly to find \(\sin(35^{\circ})\) first write it as \(\sin(35^{\circ}) - \sin(0)\)

OpenStudy (mrchair):

hi zeta XD and isn't this geometry?!?!

OpenStudy (mrchair):

zeta u think u can help with mine? is easy algebra 1 stuff and im horrible at it